# This version: August 29, 2025 )

### v3

@@ 5.1 Institutional Setting and Evidence of Strategic Behavior
We use publicly available data on Chile’s centralized college application and assignment system from 2004 to 2010 and on graduation from 2007 to 2020. The institutional setting has been described in detail by Hastings et al. (2013) and Larroucau and Rios (2020, 2021), among others. College choice in Chile is organized as a semicentralized system—a subset of universities participate in a centralized market in which a clearinghouse collects rank-order lists from applicants and determines assignments using a variant of the DA algorithm. Students can submit rank-order lists of up to eight major–university pairs (“programs”) out of more than 1,000.101010The cap was increased to ten in 2012. Priorities are program-specific and determined by a weighted average of scores obtained in a national standardized test (the PSU, for prueba de selección universitaria) and of high-school GPA. [^11] In 2014, students’ relative rank within their high school was added as one of the “primary” scores to be averaged to construct priorities. Descriptive statistics about the sample of students and programs are shown in the left panel of Tables C-1 and C-2 in Appendix C.
@@
Larroucau and Rios (2020, 2021) thoroughly document that Chilean college applicants behave strategically. Using a 2014 survey linked to administrative data on applications, they show that listed programs often do not coincide with the truly preferred programs as elicited by the survey. Focusing on applications to medicine programs, they find, for instance, that “among the 40,000 students who answered the survey, close to 10% (3,797) reported Medicine as their top preference, and 2,987 of these students ended up applying to the system. Among these, only 1,360 listed Medicine as their top preference.” Moreover, students’ probability of not including medicine at the top of their submitted list (while declaring it their most preferred program in the survey) increases as their application scores decrease, and the probability of applying to medicine drops after the 700–750 application score range, where the cutoffs for most medicine programs lie. This suggests that students tend to omit medicine as their admission chances fall, despite preferring it over other programs. These findings are confirmed in a 2019 survey, where the researchers find that “most students who did not include their first true preference in their application list expected a higher cutoff than the cutoff for their first listed preference.”
Expectations about cutoffs are a key factor behind an agent’s decision of which programs to list. Students in Chile can easily form expectations by learning from historical data on program cutoffs. In fact, many programs have relatively stable cutoffs over time (see, e.g., Kapor et al. (2024), Figure 1). Individuals know their priority scores before they decide on their submission $P$. Thus, stable cutoffs translate into a relatively high certainty about the set of schools that are feasible to the individual.
@@
Figure 1 shows that more selective programs have not only higher cutoffs but also more stable cutoffs over time. Panel (a) plots the histogram of cutoff values pooled across years and then compares that with two subsets of programs: first, with programs that we select to illustrate our methods in Section 5.2 below; and second, with medicine programs. Section 5.2 focuses on programs with a sufficiently large number of applications because we need adequate sample sizes to implement our estimation procedures. We refer to them as programs of interest. Programs of interest are more selective than most, and Medicine programs stand out as exceptionally selective. We see in Panel (a) that cutoffs generally increase with selectivity and tend to be more compressed at the top of the distribution. In order to view time variability as a share of cutoff value within each program, we divide the standard deviation of cutoffs across time by the average of cutoffs across time. Panel (b) displays the distribution of these program-specific ratios across all programs. That distribution is again compared with the programs of interest of Section 5.2 and with Medicine programs. We find that selectivity is generally associated with stability of cutoffs. For most programs of interest, the standard deviation of cutoff variability is at most 3% of cutoff value, while for medicine programs is at most 1.65%. The survey data evidence on strategic behavior regarding Medicine is consistent with agents having strong a priori knowledge about ex post cutoffs for Medicine.
@@
This figure displays distributions of cutoffs of college-program pairs in Chile for the years of 2004–2010. Histograms in both panels show distributions for three sets of programs: all programs (light gray), the subset of programs selected for the analysis of Section 5.2, i.e., programs of interest (medium gray), and medicine programs (dark gray). Panel (a) plots the cutoff values pooled over all years. Panel (b) displays the distribution of program-specific ratios: the standard deviation across the years of a program cutoff divided by the average across the years of the same program cutoff.
@@
Figure 2 provides additional evidence of strategic behavior arising from a priori knowledge of cutoffs. Consider a program $j$ with cutoff $c_{j}$. Suppose that students tend to prefer programs of higher quality, consistent with what is found in the literature. If applicants behave strategically, one would expect applications to program $j$ to peak among students with application scores close to $c_{j}$. If cutoffs tend to remain in the same neighborhood across years, students with application scores much higher than $c_{j}$ can expect to be admissible to more selective, higher-quality programs than $j$, which they prefer over $j$. Hence, we expect very few of these students to include $j$ on their list. As application scores drop and are closer to $c_{j}$, students’ chances of admission to the most selective programs decrease, and program $j$ becomes one of the most selective (desirable) programs among those for which they still have a high admission probability. Hence, we expect applications to $c_{j}$ to increase as application scores decrease and draw closer to $c_{j}$. As application scores decrease below $c_{j}$, students realize that their probability of admission to program $j$ is lower, and while program $j$ remains a relatively desirable (selective) alternative, we expect these expectations to drive applications down. This application pattern, expected if students behave strategically, is exactly what we observe in the top panel of Figure 2. Pooling all programs $j$ together, the top panel of Figure 2 shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$.
It may be difficult to disentangle the role of preferences from the role of expectations about admission probabilities when both may enter students’ choice of which programs to include in their ROLs (Manski (2004); Agarwal and Somaini (2018)). The pattern observed in the top panel of Figure 2 could, alternatively, be consistent with students not behaving strategically but preferring programs that are a good fit in terms of quality, that is, programs in which their skill level would be close to the average skill level. If this were the case, applications should peak among students whose skills (proxied by application score) are close to the mean skill level in the program. This is not what we observe in the bottom panel of Figure 2. Pooling all programs $j$ together, the bottom panel of Figure 2 shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$. Conditional on application score, the share of students applying to a program $j$ is not the highest for application scores close to the mean score among students admitted to $j$. It peaks well below this level, showing that students do not systematically prefer programs in which they would be the “average” student. This is again consistent with the hypothesis that students behave strategically.
@@
This figure provides evidence of strategic behavior in student applications. Pooling all programs $j$ together, the top panel of the figure shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$. Pooling all programs $j$ together, the bottom panel shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$.
@@ 5.2 Results
We are interested in identifying the effects of assignment to a given postsecondary program on college graduation. College returns are typically thought of as tied to college graduation, motivating our focus on graduation-related outcomes (see Kirkeboen et al. (2016) and Altonji et al. (2016) for further references). In addition, the extent to which students eventually graduate from the program to which they are assigned to (or from programs their assigned program is a pathway to) can be viewed as measure of performance of the assignment mechanism. However, for the average college program, initial enrollment and eventual graduation are far from being perfectly correlated (OECD 2019, Larroucau and Rios (2021)). We also consider graduation from a top university as another outcome of interest. The choice of this second outcome is in line with the literature on returns to education and college choice, which highlights the role of institution quality in driving returns (see Kirkeboen et al. (2016) and Altonji et al. (2016) for more references). In our setting, the set of “top” universities consists of Pontificia Universidad Católica de Chile in Santiago (PUC Santiago; hereafter PUC) and Universidad de Chile (UChile).
@@ 5.2.1 Average Structural Functions and the Importance of Preferences for Graduation Outcomes
As a simple procedure to construct a sample local to each of the admission cutoffs $c_{j}$ of interest, we use a 30-point bandwidth on either side of $c_{j}$. [^14] The resulting estimation sample is therefore invariant across outcomes once the pair $(j,k)$ is fixed. Given our choice of bandwidth, the next step of the exercise is to recover local preferences $Q_{j}$ for each student at each cutoff $c_{j}$. In the absence of strategic behavior, these preferences can be directly inferred from application lists —direct observation of $P$ and placement scores allows to compute cutoffs, budget sets, and finally $P_{j}$, which equals $Q_{j}$ under truth-telling. However, under strategic behavior, $Q_{j}$ is not fully revealed by the observed $P_{j}$. Section 4.1 shows how to construct the set of possible $Q_{j}$s, that is, the set $\bm{\mathrm{Q}}_{j}$, under various assumptions. The weak partial order assumption (hereafter WPO) implies that any submitted list is a subset of acceptable programs ordered just as in the student’s true preferences. The strong partial order assumption (hereafter SPO) adds that students who submit shorter-than-maximum lists reveal their full set of acceptable programs. Under SPO, we fully observe $Q_{j}$ for students who rank strictly fewer than the maximum number of programs (eight, in the case of our empirical application). For students ranking eight choices, $Q_{j}$ is observed only if the set $\mathbf{Q}_{j}$ comes out as a singleton. The UMAS assumption (Assumption 5, hereafter UMAS) can be used in combination with WPO and SPO to refine $\mathbf{Q}_{j}$ for the non-singleton cases (Corollary 1). Other assumptions that make sense in the empirical context of interest can be used to further refine SPO and WPO bounds. In a context with a large number of alternatives and a small cap $K$, like in Chile’s college application system, the $\bm{\mathrm{Q}}_{j}$ constructed under WPO will likely be large. As discussed in Section 4.1, additional context-specific assumptions can be combined with UMAS to reduce the size of $\bm{\mathrm{Q}}_{j}$. In the applied literature, fields of study are widely recognized as key drivers of student preferences—and of the heterogeneity in preferences across students—for post-secondary programs (Altonji et al., 2016). Building on this insight, we propose the following fields assumption:
@@
$\forall\mathcal{F},P(\omega)\cap\mathcal{F}=\emptyset\Rightarrow Q_{j}(\omega)\notin\mathcal{F}\times\mathcal{F}.$
@@
The figure shows estimated bounds for $\mathbb{E}[Y(k)|Q_{j}=(j,k),S_{j}=c_{j}]$ (y-axis) with $k=$ Bachillerato de Ingreso Comun at UChile for several local first-best programs $j$ (x-axis), two binary graduation outcomes (left and right panels), and four different behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each $j$ in each panel). The outcome in the right panel is graduation from a top university (PUC or UChile) and graduation from the local first-best choice $j$ in the left panel. The four behavioral assumptions are, from left to right, SPO, WPO in combination with UMAS and fields, WPO in combination with UMAS, and WPO alone. The average parameter denotes the probability of graduating (from a top university or from program $j$) among those individuals with $S_{j}=c_{j}$ and $Q_{j}=(j,k)$ if they were all assigned to program $k$. All sets of bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff $c_{j}$ of the program of interest. Note that these are estimated bounds and not inference bounds.
@@
We begin by discussing results under WPO with UMAS and the fields assumption, which we take as our preferred specification. This combination retains the minimal behavioral restrictions of WPO while incorporating two intuitive and contextually grounded assumptions—UMAS and fields—that meaningfully refine the set of plausible preferences and yield relatively tight bounds. For both outcomes, we find heterogeneous effects across different local first-best programs. This is evident from the non-overlapping bounds across different $j$. For instance, in the left panel, students whose local first-best is Medicine at UChile are more likely to graduate from that program after assignment to the Bachillerato than those whose first-best is Medicine at Universidad de Santiago de Chile or Math at PUC. In the right panel, students with Medicine at UChile as their first-best are more likely to graduate from a top university when assigned to the Bachillerato than those with Math at PUC as their first-best.
@@
We therefore interpret the results in Figure 3 as evidence that students’ preferences matter for their graduation outcomes, suggesting that preferences are correlated with effort choices or ability that is not captured by test scores. [^16] Evidence from Heckman and Kautz (2012) shows that achievement tests do not fully capture soft skills or personality traits that influence labor market outcomes (see also Borghans et al. (2008), Almlund et al. (2011)). Structural models support this, showing that factors beyond test scores affect educational choices. For example, Arcidiacono (2005) find that, even conditional on SAT scores, some individuals have higher college admission chances, financial aid likelihood, labor market earnings without college, and returns to all majors.
@@
We next examine how different behavioral assumptions affect the estimated bounds. Table C-4 in Appendix provides descriptive statistics and the estimated $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ for each pair $(j,k)$ and each set of assumptions. The different assumptions are used to construct $\bm{\mathrm{Q}}_{j}$ for each individual and therefore determine the conditioning set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and the values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$. Under WPO, $\bm{\mathrm{Q}}_{j}$ is larger than under SPO, resulting in a larger conditioning set (as shown in Table C-4, Columns (1) vs. (4)). Given a fixed bandwidth around the cutoff $c_{j}$, the set $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ being larger under WPO than under SPO means that the denominator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is larger under WPO than under SPO. In addition, as the set $\bm{\mathrm{Q}}_{j}$ is larger when constructed under WPO than when constructed under SPO, the share of individuals for whom $Q_{j}$ is found to be a singleton is lower under WPO than it is under SPO. This means that the numerator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is smaller under WPO than under SPO. Overall, this leads to $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ being smaller under WPO than under SPO, as seen in Table C-4. The increase in $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ as we impose SPO instead of WPO unambiguously tends to reduce the width of the bounds. The exact extent to which the width of the bounds shrinks as we impose SPO instead of WPO also depends on the joint distribution of the outcome $Y$ and $\bm{\mathrm{Q}}_{j}$, however. Imposing UMAS eliminates certain elements from individuals’ $\bm{\mathrm{Q}}_{j}$s. As a consequence, following the same logic as above, imposing UMAS shrinks the conditioning set $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ and increases $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ (see Columns (3) vs. (4)). Again, the increased values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ are the key force that tends to reduce the width of the bounds as UMAS is imposed. The fields assumption, when added, further trims the set $\bm{\mathrm{Q}}_{j}$. Students that are part of $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ without the fields assumption but do not report programs from $\mathcal{F}(j)$ or $\mathcal{F}(k)$ are not part of $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ with the fields assumption. As before, this increases $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ (see Columns (2) vs. (3)).
Note, however, that the impact of the fields assumption depends on the overlap between fields of $j$ and $k$. Under stability, the first coordinate of $P_{j}$ and the first coordinate of any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just above $c_{j}$, while the second coordinates of $P_{j}$ and any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just below $c_{j}$ (Proposition 2). This means that, absent the fields assumption, each student in the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ includes a program from either $\mathcal{F}(j)$ or $\mathcal{F}(k)$ in her reported preference. As a consequence, the fields assumption can only reduce the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ for pairs $(j,k)$ for which $\mathcal{F}(j)\neq\mathcal{F}(k)$. In our empirical context, the Bachillerato de Ingreso Comun belongs to the STEM field. Therefore, the fields does not affect results for the pair with Math at PUC as $j$, since both $j$ and $k$ fall within STEM. However, for pairs where $j$ belongs to Health/Medicine, the assumption does help reduce the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and narrow the bounds.

### v4

@@ 5.1 Institutional Setting and Evidence of Strategic Behavior
We use publicly available data on Chile’s centralized college application and assignment system from 2004 to 2010 and on graduation from 2007 to 2020. The institutional setting has been described in detail by hastingsneilsonzimmerman2013 and larroucaurios2020; larroucaurios2021, among others. College choice in Chile is organized as a semicentralized system—a subset of universities participate in a centralized market in which a clearinghouse collects rank-order lists from applicants and determines assignments using a variant of the DA algorithm. Students can submit rank-order lists of up to eight major–university pairs (“programs”) out of more than 1,000.121212The cap was increased to ten in 2012. Priorities are program-specific and determined by a weighted average of scores obtained in a national standardized test (the PSU, for prueba de selección universitaria) and of high-school GPA. [^13] In 2014, students’ relative rank within their high school was added as one of the “primary” scores to be averaged to construct priorities. Descriptive statistics about the sample of students and programs are shown in Tables C-1 and C-2.
@@
larroucaurios2020; larroucaurios2021 thoroughly document that Chilean college applicants behave strategically. Using a 2014 survey linked to administrative data on applications, they show that listed programs often do not coincide with the truly preferred programs as elicited by the survey. Focusing on medicine programs, they show that as application scores decline, students become more likely to exclude medicine from the top of their submitted list—even when ranking it first in the survey—and application rates drop sharply in the 700–750 range, where most medicine cutoffs lie. This pattern indicates that students forgo medicine as their admission prospects weaken, despite preferring it over other programs.
@@
Expectations about cutoffs play a central role in program choice. In Chile, students can readily infer these from historical cutoff data, and many programs exhibit stable cutoffs over time (see, e.g., kapor2024, Figure 1). Because students observe their priority scores before submitting $P$, cutoff stability implies considerable certainty about which programs are feasible. Figure 1 shows that more selective programs have not only higher cutoffs but also more predictable cutoffs over time. Panel (a) plots the histogram of cutoff values pooled across years and then compares that with two subsets of programs: first, with programs that we select to illustrate our methods in Section 5.2 below; and second, with medicine programs. Section 5.2 focuses on programs with a sufficiently large number of applications because we need adequate sample sizes to implement our estimation procedures. We refer to them as programs of interest. Programs of interest are more selective than most, and Medicine programs stand out as exceptionally selective. We see in Panel (a) that cutoffs generally increase with selectivity and tend to be more compressed at the top of the distribution. In order to assess how predictable program-specific cutoffs are, we estimate an auto-regressive (AR) model for cutoff value at time $t$ for each program as a function of past cutoff values. Our sample has seven years of data which allows us to estimate an AR with three lags after de-meaning the data. Panel (b) displays the distribution of these program-specific $R^{2}$’s. That distribution is again compared with the programs of interest of Section 5.2 and with medicine programs. We find that selectivity is generally associated with predictability of cutoffs. For most programs of interest, the $R^{2}$ is larger than 0.70, while 70% of medicine programs have an $R^{2}$ greater than 0.90. The survey data evidence on strategic behavior regarding Medicine is consistent with agents having strong a priori knowledge about ex post cutoffs for Medicine.
@@
This figure displays distributions of cutoffs of college-program pairs in Chile for the years of 2004–2010. Histograms in both panels show distributions for three sets of programs: all programs (light gray), the subset of programs selected for the analysis of Section 5.2, i.e., programs of interest (medium gray), and medicine programs (dark gray). Panel (a) plots the cutoff values pooled over all years. Panel (b) displays the distribution of program-specific $R^{2}$’s. We construct an $R^{2}$ for each program by regressing cutoff at time $t$ on three lagged cutoff values after de-meaning the seven years of cutoff data.
@@
Figure 2 provides additional evidence of strategic behavior arising from a priori knowledge of cutoffs. Consider a program $j$ with cutoff $c_{j}$. Suppose that students tend to prefer programs of higher quality, consistent with what is found in the literature. If applicants behave strategically, one would expect applications to program $j$ to peak among students with application scores close to $c_{j}$. If cutoffs tend to remain in the same neighborhood across years, students with application scores much higher than $c_{j}$ can expect to be admissible to more selective, higher-quality programs than $j$, which they prefer over $j$. Hence, we expect very few of these students to include $j$ on their list. As application scores drop and are closer to $c_{j}$, students’ chances of admission to the most selective programs decrease, and program $j$ becomes one of the most selective (desirable) programs among those for which they still have a high admission probability. Hence, we expect applications to $c_{j}$ to increase as application scores decrease and draw closer to $c_{j}$. As application scores decrease below $c_{j}$, students realize that their probability of admission to program $j$ is lower, and while program $j$ remains a relatively desirable (selective) alternative, we expect these expectations to drive applications down. This application pattern, expected if students behave strategically, is exactly what we observe in the top panel of Figure 2. Pooling all programs $j$ together, Figure 2 shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$.
It may be difficult to disentangle the role of preferences from the role of expectations about admission probabilities when both may enter students’ choice of which programs to include in their ROLs (manski_ecta2004; agarwal2018). The pattern observed in Figure 2 could, alternatively, be consistent with students not behaving strategically but preferring programs that are a good fit in terms of quality, that is, programs in which their skill level would be close to the marginal skill level. If this were the case, application behavior would not change discontinuously as students’ scores cross a program’s admission cutoff. Although such discontinuities are not required for strategic behavior to arise, examining specific program pairs—rather than pooling all programs as in Figure 2—reveals clear discontinuities (Figure 3). The left panels plot, for three program pairs $(j,k)$, the density of the application score $S_{j}$ among applicants with $P_{j}=(j,k)$ and show a significant jump at $S_{j}=c_{j}$. The right panels display the probability that $P_{j}=(j,k)$ among all potential applicants as a function of $S_{j}$, again exhibiting a sharp discontinuity at the cutoff. Importantly, the programs $j$ in these pairs have highly predictable cutoffs: the $R^{2}$ from three-lag autoregressive models forecasting the cutoff at time $t$ exceeds 0.98 in all three cases. This discontinuous type of strategic behavior may lead to issues with internal validity of an RD identification strategy that controls for $P_{j}$ (see Section 5.2.3 for discussion).
@@
This figure provides evidence of strategic behavior in student applications. Pooling all programs $j$ together, the figure shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$.
Figure 3: Evidence of strategic behavior of the discontinuous type
This figure provides evidence of strategic behavior of the discontinuous type. Plots on the left show, for three pairs $(j,k)$, the density of the application score $S_{j}$ among applicants for whom $P_{j}=(j,k)$. Plots were generated using the rddensity package; the discontinuity estimates provided were obtained using the default MSE-based bandwidth selection method. The graphs on the right show the probability that $P_{j}=(j,k)$ among all potential applicants as a function of $S_{j}$. The discontinuity estimates provided were obtained using the rdrobust package and its default MSE-based bandwidth selection method.
@@ 5.2 Results
We are interested in identifying the effects of assignment to a given postsecondary program on college graduation. College returns are typically thought of as tied to college graduation, motivating our focus on graduation-related outcomes (see kirkeboen2016 and altonjiarcidiaconomaurel2016 for further references). In addition, the extent to which students eventually graduate from the program to which they are assigned to (or from programs their assigned program is a pathway to) can be viewed as measure of performance of the assignment mechanism. However, for the average college program, initial enrollment and eventual graduation are far from being perfectly correlated (OECD 2019, larroucaurios2021). We also consider graduation from a top university as another outcome of interest. The choice of this second outcome is in line with the literature on returns to education and college choice, which highlights the role of institution quality in driving returns (see kirkeboen2016 and altonjiarcidiaconomaurel2016 for more references). In our setting, the set of “top” universities consists of Pontificia Universidad Católica de Chile in Santiago (PUC Santiago; hereafter PUC) and Universidad de Chile (UChile).
@@ 5.2.1 Average Structural Functions and the Importance of Preferences for Graduation Outcomes
As a simple procedure to construct a sample local to each of the admission cutoffs $c_{j}$ of interest, we use a 30-point bandwidth on either side of $c_{j}$. [^16] The resulting estimation sample is therefore invariant across outcomes once the pair $(j,k)$ is fixed. Given our choice of bandwidth, the next step of the exercise is to recover local preferences $Q_{j}$ for each student at each cutoff $c_{j}$. In the absence of strategic behavior, these preferences can be directly inferred from application lists —direct observation of $P$ and placement scores allows to compute cutoffs, budget sets, and finally $P_{j}$, which equals $Q_{j}$ under truth-telling. However, under strategic behavior, $Q_{j}$ is not fully revealed by the observed $P_{j}$. Section 4.1 shows how to construct the set of possible $Q_{j}$s, that is, the set $\bm{\mathrm{Q}}_{j}$, under various assumptions. The weak partial order assumption (WPO) implies that any submitted list is a subset of acceptable programs ordered just as in the student’s true preferences. The strong partial order assumption (SPO) adds that students who submit shorter-than-maximum lists reveal their full set of acceptable programs. Under SPO, we fully observe $Q_{j}$ for students who rank strictly fewer than the maximum number of programs (eight, in the case of our empirical application). For students ranking eight choices, $Q_{j}$ is observed only if the set $\mathbf{Q}_{j}$ comes out as a singleton. The UMAS assumption (Assumption 5, hereafter UMAS) can be used in combination with WPO and SPO to refine $\mathbf{Q}_{j}$ for the non-singleton cases (Corollary 1). In a context with a large number of alternatives and a small cap $K$, like in Chile’s college application system, the $\bm{\mathrm{Q}}_{j}$ constructed under WPO will likely be large. As discussed in Section 4.1, additional context-specific assumptions can be combined with UMAS to reduce the size of $\bm{\mathrm{Q}}_{j}$. In the applied literature, fields of study are widely recognized as key drivers of student preferences—and of the heterogeneity in preferences across students—for post-secondary programs (altonjiarcidiaconomaurel2016). Building on this insight, we propose the following fields assumption:
@@
$\forall\mathcal{F},P(\omega)\cap\mathcal{F}=\emptyset\Rightarrow Q_{j}(\omega)=(a,b):a\notin\mathcal{F}\text{ and }b\notin\mathcal{F}.$
@@
The figure shows estimated bounds for $\mathbb{E}[Y(k)|Q_{j}=(j,k),S_{j}=c_{j}]$ (y-axis) with $k=$ Bachillerato de Ingreso Comun at UChile for several local first-best programs $j$ (x-axis), two binary graduation outcomes (graduation from the local first-best choice $j$ on the left panel and graduation from a top university, i.e., PUC or UChile, on the right panel), and four different behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each $j$ in each panel). The four behavioral assumptions are, from left to right, SPO+UMAS+fields, SPO+UMAS, WPO+UMAS+fields, and WPO+UMAS. The average parameter denotes the probability of graduating (from a top university or from program $j$) among those individuals with $S_{j}=c_{j}$ and $Q_{j}=(j,k)$ if they were all assigned to program $k$. All sets of bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff $c_{j}$ of the program of interest. Note that these are estimated bounds and not inference bounds.
@@
For both outcomes, we find evidence of heterogeneous effects across different local first-best programs —in a number of cases, bounds do not overlap across $j$. For instance, in the left panel, students whose local first-best is Medicine at UChile are more likely to graduate from that program after assignment to the Bachillerato than those whose first-best is Medicine at Universidad de Santiago de Chile or Math at PUC. In the right panel, students with Medicine at UChile as their first-best are more likely to graduate from a top university when assigned to the Bachillerato than those with Math at PUC as their first-best.
@@
We therefore interpret the results in Figure 4 as evidence that students’ preferences matter for their graduation outcomes, suggesting that preferences are correlated with effort choices or ability that is not captured by test scores. [^18] Evidence from heckmankautz2012 shows that achievement tests do not fully capture soft skills or personality traits that influence labor market outcomes (see also borghans2008, almlund2011). Structural models support this, showing that factors beyond test scores affect educational choices. For example, arci2005 find that, even conditional on SAT scores, some individuals have higher college admission chances, financial aid likelihood, labor market earnings without college, and returns to all majors.
@@
We next examine how different behavioral assumptions affect the estimated bounds. Table C-4 in the appendix provides descriptive statistics and the estimated $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ for each pair $(j,k)$ and each set of assumptions. The different assumptions are used to construct $\bm{\mathrm{Q}}_{j}$ for each individual and therefore determine the conditioning set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and the values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$. Under WPO, $\bm{\mathrm{Q}}_{j}$ is larger than under SPO, resulting in a larger conditioning set (Table C-4, Columns (1) vs. (3) and (2) vs. (4)). Given a fixed bandwidth around the cutoff $c_{j}$, this means that (i) the denominator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is larger under WPO than under SPO; and (ii) the share of individuals for whom $Q_{j}$ is found to be a singleton is lower under WPO than it is under SPO, which in turn implies that the numerator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is smaller under WPO than under SPO. Overall, this leads to $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ being smaller under WPO than under SPO. The increase in $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ as we impose SPO instead of WPO unambiguously tends to reduce the width of the bounds. The exact extent to which the width of the bounds shrinks as we impose SPO instead of WPO also depends on the joint distribution of the outcome $Y$ and $\bm{\mathrm{Q}}_{j}$, however. Subsequently imposing UMAS and the fields assumption eliminate certain elements from individuals’ $\bm{\mathrm{Q}}_{j}$s and futher increase $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ (Columns (1) vs. (2) and (3) vs. (4)).  [^19] Note that the impact of the fields assumption depends on the overlap between fields of $j$ and $k$. Under stability, the first coordinate of $P_{j}$ and the first coordinate of any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just above $c_{j}$, while the second coordinates of $P_{j}$ and any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just below $c_{j}$ (Proposition 2). This means that, absent the fields assumption, each student in the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ includes a program from either $\mathcal{F}(j)$ or $\mathcal{F}(k)$ in her reported preference. As a consequence, the fields assumption can only reduce the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ for pairs $(j,k)$ for which $\mathcal{F}(j)\neq\mathcal{F}(k)$.